OpenStudy (grandmoffchris):

3. At a bargain store, Tanya bought 3 items that each cost the same amount. Tony bought 4 items that each cost the same amount, but each was $2.25 less than the items that Tanya bought. Both Tanya and Tony paid the same amount of money. What was the individual cost of each person's items? (a) Write an equation. Let x represent the cost of one of Tanya's items. (b) Solve the equation. Show your work. (c) Check your solution. Show your work. (d) State the solution in complete sentences.

OpenStudy (grandmoffchris):

for A its 3x = 4x - 2.25?

OpenStudy (studygurl14):

Almost. It would be 3x = 4(x-2.25)

OpenStudy (grandmoffchris):

Oh, for B and C and D I will need help

OpenStudy (studygurl14):

Okay, so solve the equation. First step is to distribute the 4

OpenStudy (grandmoffchris):

distribute?

OpenStudy (grandmoffchris):

wait do I rest it from both numbers or do I add it?

OpenStudy (studygurl14):

3x = 4(x - 2.25) --> 3x = 4x + 4(-2.25) = ?

OpenStudy (grandmoffchris):

4 x -2.25?

OpenStudy (studygurl14):

What is 4 times -2.25?

OpenStudy (grandmoffchris):

9

OpenStudy (grandmoffchris):

its 9 right?

OpenStudy (studygurl14):

almost. -9 So you get 3x = 4x - 9

OpenStudy (studygurl14):

Now subtract 4x from both sides

OpenStudy (grandmoffchris):

-1x and -5?

OpenStudy (grandmoffchris):

?

OpenStudy (studygurl14):

almost -x = -9 So x = 9

OpenStudy (grandmoffchris):

@StudyGurl14

OpenStudy (grandmoffchris):

Im back, sorry I had to do something

OpenStudy (grandmoffchris):

@StudyGurl14

OpenStudy (austin.l):

Okay, so now that you have x=9 plug that back into your original equation to see if it holds true. \[3x = 4(x-2.25)\] \[3(9) = 4((9)-2.25)\] Are those equal?

OpenStudy (grandmoffchris):

No

OpenStudy (grandmoffchris):

wait yes

OpenStudy (grandmoffchris):

@Austin.L

OpenStudy (grandmoffchris):

@inkyvoyd

OpenStudy (austin.l):

They should be lol

OpenStudy (austin.l):

The last part, you literally type out what you did in english sentences.

OpenStudy (grandmoffchris):

So the answer for C is what..?

OpenStudy (austin.l):

Do the math for what I posted.

OpenStudy (grandmoffchris):

33.75?

OpenStudy (austin.l):

What is 3*9?

OpenStudy (grandmoffchris):

27

OpenStudy (grandmoffchris):

4 x 9 = 36

OpenStudy (austin.l):

Right, Now 4*2.25=9 36-9=27 Meaning they are equal.

OpenStudy (grandmoffchris):

Oh okay, and so thats the answer?

OpenStudy (austin.l):

For C. Jus the work for it.

OpenStudy (grandmoffchris):

Ok

OpenStudy (grandmoffchris):

thank you so much